OpenStudy (zned6559):

Please help I will medal why does d/dx S from 0 to pi cosx^2 dx=0

OpenStudy (zned6559):

\[\frac{ d }{ dx }\int\limits_{0}^{\pi}cosx^2dx = 0\]

OpenStudy (zned6559):

why does it equal zero???

OpenStudy (michele_laino):

since the integral is a real number

OpenStudy (michele_laino):

and not a function

OpenStudy (zned6559):

i still don't get it

OpenStudy (zned6559):

what does it mean when d/dx is on the outside versus being on the inside of the integral?

OpenStudy (michele_laino):

we can write this: \[\frac{d}{{dx}}\int_0^\pi {{{\left( {\cos x} \right)}^2}} dx = \frac{d}{{dx}}\left. {\left( {\frac{{x + \sin x\cos x}}{2}} \right)} \right|_0^\pi = \frac{d}{{dx}}\left( {\frac{\pi }{2}} \right) = 0\]

OpenStudy (zned6559):

how did you get the stuff inside the parenthesis?

OpenStudy (michele_laino):

I have used this identity: \[{\left( {\cos x} \right)^2} = \frac{{1 + \cos \left( {2x} \right)}}{2}\]

ganeshie8 (ganeshie8):

If we pretend that d/dx isn't there, would you agree that definite integral evaluates to some real number ?

OpenStudy (zned6559):

yes

ganeshie8 (ganeshie8):

whats the derivative of a real number ?

OpenStudy (zned6559):

0

OpenStudy (zned6559):

so what if the d/dx is on the inside?

ganeshie8 (ganeshie8):

it isn't on the inside

ganeshie8 (ganeshie8):

it is outside to make your life easy :)

ganeshie8 (ganeshie8):

if it is inside, you will have to find the derivative first and then evaluate the integral